OpenStudy (anonymous):

Please Help:(( Input the slope and y -intercept without solving for y . Reduce all fractional answers to lowest terms. 2x + 3y = 6 m= b=

OpenStudy (anonymous):

What does m and b mean

ganeshie8 (ganeshie8):

Ax + By = C m = -A/B b = C/B